OpenStudy (anonymous):

Please answer i nedd this tomorrow? a slow jogger runs a mile in 13 minutes? What is the speed in a. in/sec b. Km.hr?

OpenStudy (abhisar):

Hello @jella ! \(\boxed{\boxed{\boxed{\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\color{blue}{\mathfrak{~~~~Welcome~to~OpenStudy!~\ddot\smile}}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}}}}\)

OpenStudy (abhisar):

1 mile = 63360 inches and 13 minutes = 13*60 = 780, so jogger speed will be \[\frac{ 63360 }{780} = 81.23~inch/\sec\]

OpenStudy (abhisar):

Now again 1 mile = 1.6 km and 13 min = 13*0.016 = 0.2\[\frac{ 1.6 }{0.2} = 8km/hr\]hr, so joggers speed will be